OpenStudy (anonymous):

If the volume of a sphere is 36π cubic units, what is the radius?

OpenStudy (unklerhaukus):

\[V_{\text{sphere}}=\frac 43\pi r^3\]

OpenStudy (unklerhaukus):

solve for the \(r\)adius

OpenStudy (anonymous):

okay thanks is it 3?

OpenStudy (unklerhaukus):

\[\frac 43\pi\times 3^3=36\pi=V\]\[r=3\checkmark \]

OpenStudy (anonymous):

thanks